The positive integer just greater than (1+0.0001)^10000

Grade:11

2 Answers

Aditya
7 years ago
Using binomial expansion and |x|
DR STRANGE
7 years ago
It is well known that lim n—> inf (1 + 1/n)^n evaluates to e.
So if we put n = 10000
(1 + 1/10000)^10000
And by binomial theorem,
(1 + 1/10000)^10000 = 1 + 10000C1 1/1000 + … > 2
Hence the answer is 3
please approve if you are satisfied n correct me if I am wrong